# Prompt-Specific Poisoning Attacks on Text-to-Image Generative Models

### v2

@@ Abstract
Trained on billions of images, diffusion-based text-to-image models seem impervious to traditional data poisoning attacks, which typically require poison samples approaching 20% of the training set. In this paper, we demonstrate the surprising result that state-of-the-art text-to-image generative models are in fact highly vulnerable to poisoning attacks. Our work is driven by two key insights. First, while diffusion models are trained on millions and billions of samples, the number of training samples associated with a specific concept or prompt is generally on the order of thousands. This suggests that these models will be vulnerable to prompt-specific poisoning attacks that corrupt a model’s ability to respond to specific targeted prompts. Second, poison samples can be carefully crafted to maximize poison potency to ensure success with very few samples.
@@ 1 Introduction
Over the last year, diffusion based text-to-image models have taken the Internet by storm, growing from research projects to numerous applications in advertising, fashion [1, 2], web development [3, 4, 5], and AI art [6, 7, 8, 9]. Models like Stable Diffusion SDXL, Midjourney v5, Dalle-3, Imagen, Adobe Firefly and others boast tens of millions of registered users and have produced billions of images [10].
Despite their significant and disruptive impact on business and creative industries, few have considered the vulnerability of diffusion models to data poisoning attacks. Poisoning attacks manipulate training data to introduce unexpected behavior to the model at training time. They have been studied extensively in the context of classification tasks using deep neural networks (DNN). Poisoning attacks cause predictable misclassifications, but typically demand a substantial volume of poison data for success, e.g., ratio of poison training samples to benign samples of 20% or higher. Since today’s diffusion models are trained on hundreds of millions (or billions) of images, a common assumption is that poisoning attacks on these models would require millions of poison samples, making them infeasible in practice.
@@ 2.2 Data Poisoning Attacks
Defenses against data poisoning are also well studied. Some [46, 47, 48, 49, 50] focus on detecting poison data by leveraging their unique behavior while others [51, 52, 53] advocate for robust training to mitigate the influence of poison data during training time. However, poison defenses continue to face challenges, particularly as more potent, adaptive attacks frequently find ways to bypass existing defenses [40, 54, 55, 56, 57].
@@ 3 Feasibility of Poisoning Diffusion Models
In this work, we demonstrate the unexpected finding that generic text-to-image diffusion models, despite having massive training datasets, are susceptible to data poisoning attacks. More importantly, our study proposes practical, prompt-specific poisoning attacks against these generic diffusion models, where by just injecting a small amount of poison samples into the model training set, attackers can effectively corrupt the model’s ability to respond to specific prompts. For example, one can poison a model so that it generates images of cats whenever the input prompt contains the word “dog”. Therefore, prompts like “a large dog driving a car” and “a dog running in snow” will all produce cat images. Figure 1 illustrates the high-level attack process. Note that our attacks do not require modifications to the model training pipeline or the diffusion process, in constrast with existing attacks discussed in §2.
@@ 3.1 Threat Model
We note that unlike prior works on poisoning text-to-image diffusion models (§2), our attack does not require priviledged access to the model training and deployment. Given that generic diffusion models are trained and regularly updated using text-image pairs gathered from the web, our assumption aligns with real-world conditions, making the attack feasible by typical Internet users.
@@ 3.2 Concept Sparsity Induces Vulnerability
Existing research finds that an attack must poison a decent percentage of the model’s training dataset to be effective. For DNN classifiers, the poisoning ratio should exceed $5$% for backdoor attacks [39, 62] and $20$% for indiscriminative attacks [63, 64]. A recent backdoor attack against diffusion models needs to poison half of the dataset [13]. Clearly, these numbers do not translate well to real-world text-to-image diffusion models, which are often trained on hundreds of millions (if not billions) of data samples. Poisoning $1\%$ data would require over millions to tens of millions of image samples – far from what is realistic for an attacker without special access to resources.
@@ 6.1 Experimental Setup
Nightshade Attack Configuration.   Following the attack design in §5.3, we randomly select $5$K samples from LAION-5B (minus LAION-Aesthetic) as the natural dataset $\{\mathsf{Text/Image}\}$. We ensure they do not overlap with the 100K training samples in Table II. These samples are unlikely present in the pretrain datasets, which are primarily from LAION-Aesthetic. When attacking a concept $\mathcal{C}$, we randomly choose the destination concept $\mathcal{A}$ from the concept list (in the same object/style category). For guided perturbation, we follow prior work to use LPIPS budget of $p=0.07$ and run an Adam optimizer for 500 steps [78, 14]. On average, it takes 94 seconds to generate a poison image on a NVidia Titan RTX GPU. Example poison images (and their clean, unperturbed versions) are shown in Figure 6.
@@ 6.2 Attack Effectiveness
Visualizing changes in model internals.   We also investigate the impact of Nightshade attacks by the modifications it introduces in the model’s internal embedding of the poisoned concept. Specifically, we study the cross-attention layers, which encode the relationships between text tokens and a given image [83, 81]. Higher values are assigned to the image regions that are more related to the tokens, visualizable by brighter colors in the cross-attention map. Figure 13 plots the cross-attention maps of a model before and after poisoning model (SD-V2 with $200$ poison data) for two object concepts targeted by Nightshade (“hat” and “handbag”). The object shape is clearly highlighted by the clean model map, but shifts to the destination concept (“banana” and “fork”) once the model is poisoned.
@@ 6.3 Impact of Clean Training Data
Subsequent continous training on clean data only.   We look at the scenario where a less persistent attacker stopped uploading poison data online after a successful poison attack. Over time, the poison effect may decrease as model trainer continously updates the poisoned model on only clean data. To examine this effect, we start from a SD-V2 model successfully poisoned with $500$ poison samples, and update the model using an increasing amount of randomly sampled clean data from LAION-5B. Figure 19 in the Appendix shows that the attack success rate does decrease with the # of new clean data. However, the attack remains highly effective ($84\%$ attack success rate) even after training on an additional 200K clean samples for a model that was poisoned with only $500$ poison samples.
@@ 6.4 Bleed-through to Other Concepts
6.5 Stacking Multiple Nightshade Attacks
@@
Figure 17 illustrates the impact of these attacks with example images generated on prompts not targeted by any poison attacks. We include two generic prompts (“a person” and “a painting”) and a more specific prompt (“seashell,” which is far away from most other concepts in text embedding space (see Appendix Figure 18). Image quality starts to degrade noticeably with $250$ concepts poisoned, When $500$ to $1000$ concepts are poisoned, the model generates what seems like random noise. For a model training from scratch (LD-CC), similar levels of degradation requires $500$ concepts to be poisoned (Table IX in Appendix). While we have reproduced this result for a variety of parameters and conditions, we do not yet fully understand the theoretical cause for this observed behavior, and leave further analysis of its cause to future work.
@@ 6.6 Attack Generalizability
Attack performance on diverse prompts.   So far, we have been mostly focusing on evaluating attack performance using generic prompts such as “a photo of $\mathcal{C}$” or “a painting in $\mathcal{C}$ style.” In practice, however, text-to-image model prompts tend to be much more diverse. Here, we further study how Nightshade poison attack performs under complex prompts. Given a poisoned concept $\mathcal{C}$, we follow prior work [37] to generate $4$ types of complex prompts (examples shown in Table VI). More details on the prompt construction can be found in Section 4 of [37]. We summarize our results in Table VI. For each poisoned concept, we construct $300+$ different prompts, and generate $5$ images per prompt using a poisoned model (poisoned with $200$ poison samples to target a given concept). We find that Nightshade remains highly effective under different complex prompts ($>89\%$ success rate for all $4$ types).
@@ 7 Potential Defenses
Automated image captioning.   Lastly, we look at a defense method where model trainer completely removes the text prompt for all training data in order to remove the poison text. Once removed, model trainer can leverage existing image captioning tools [93, 94] to generate new text prompts for each training image. Similar approaches have been used to improve the data quality of poorly captioned images [95, 96].
@@ 8 Poison Attacks as Copyright Protection
While we have not yet finalized a public release of Nightshade, we have been approached by and are in discussions with several companies in the music, photography and gaming industries to deploy Nightshade on their copyrighted content. The datasets in question include several hundred million images. Finally, relevant model companies Google, Meta, Stability.ai and OpenAI have all been made aware of this work prior to this submission.
@@ A.3 Additional Results of Simple Dirty-Label Poisoning Attacks
A.4 Additional Results on Bleed through and Stacking Multiple Attacks
We evaluate the “related” concept bleed-through effects between artists and the art styles they are known for. We include $195$ artists associated with $28$ styles from the Wikiart dataset [72]. We poison each art style $\mathcal{C}$, then test poison’s impact on generating painting of artists whose style belong to style $\mathcal{C}$, without mentioning the poisoned style $\mathcal{C}$ in the prompt, e.g., query with “a painting by Picasso” for models with “cubism” poisoned. Table VII shows that with $200$ poison data on art style, Nightshade achieves $>91\%$ CLIP attack success rate on artist names alone, similar to its performance on the poisoned art style.
@@ A.4 Additional Results on Bleed through and Stacking Multiple Attacks
Figure 19: Nightshade’s attack success rate (CLIP-based) decreases when mdoel trainer continously trains an already-poisoned model on an increasing number of clean data. The base model is poisoned with 100, 300, and 500 poison data.

### v3

# Nightshade: Prompt-Specific Poisoning Attacks on Text-to-Image Generative Models

@@ Abstract
Trained on billions of images, diffusion-based text-to-image models seem impervious to traditional data poisoning attacks, which typically require poison samples approaching 20% of the training set. In this paper, we show that state-of-the-art text-to-image generative models are in fact highly vulnerable to poisoning attacks. Our work is driven by two key insights. First, while diffusion models are trained on billions of samples, the number of training samples associated with a specific concept or prompt is generally on the order of thousands. This suggests that these models will be vulnerable to prompt-specific poisoning attacks that corrupt a model’s ability to respond to specific targeted prompts. Second, poison samples can be carefully crafted to maximize poison potency to ensure success with very few samples.
@@ 1 Introduction
Since 2022, diffusion based text-to-image models have taken the Internet by storm, growing from research projects to numerous applications in advertising, fashion [1, 2], web development [3, 4, 5], and AI art [6, 7, 8, 9]. Models like Stable Diffusion SDXL, Midjourney v5, Dalle-3, Imagen, Adobe Firefly and others boast tens of millions of registered users and have produced billions of images [10].
To date, public consensus considers these diffusion models impervious to data poisoning attacks. Poisoning attacks manipulate training data to introduce unexpected behavior to the model at training time, and have been studied extensively in the context of classification tasks using deep neural networks (DNN). Poisoning attacks cause predictable misclassifications, but typically demand a substantial volume of poison data for success, e.g., ratio of poison training samples to benign samples of 20% or higher. Since today’s diffusion models are trained on hundreds of millions (or billions) of images, a common assumption is that poisoning attacks on these models would require millions of poison samples, making them infeasible in practice.
@@ 2.2 Data Poisoning Attacks
Defenses against data poisoning are also well studied. Some [46, 47, 48, 49, 50, 51] focus on detecting poison data by leveraging their unique behavior while others [52, 53, 54] advocate for robust training to mitigate the influence of poison data during training time. However, poison defenses continue to face challenges, particularly as more potent, adaptive attacks frequently find ways to bypass existing defenses [40, 55, 56, 57, 58].
@@ 3 Feasibility of Poisoning Diffusion Models
In this work, we demonstrate the unexpected finding that generic text-to-image diffusion models, despite having massive training datasets, are susceptible to data poisoning attacks. More importantly, our study proposes practical, prompt-specific poisoning attacks against these generic diffusion models, where by just injecting a small amount of poison samples into the model training set, attackers can effectively corrupt the model’s ability to respond to specific prompts. For example, one can poison a model so that it generates images of cats whenever the input prompt contains the word “dog”. Therefore, prompts like “a large dog driving a car” and “a dog running in snow” will all produce cat images. Figure 1 illustrates the high-level attack process. Note that our attacks do not require modifications to the model training pipeline or the diffusion process, in contrast with existing attacks discussed in §2.
@@ 3.1 Threat Model
We note that unlike prior works on poisoning text-to-image diffusion models (§2), our attack does not require privileged access to the model training and deployment. Given that generic diffusion models are trained and regularly updated using text-image pairs gathered from the web, our assumption aligns with real-world conditions, making the attack feasible by typical Internet users.
@@ 3.2 Concept Sparsity Induces Vulnerability
Existing research finds that an attack must poison a decent percentage of the model’s training dataset to be effective. For DNN classifiers, the poisoning ratio should exceed $5$% for backdoor attacks [39, 63] and $20$% for indiscriminate attacks [64, 65]. A recent backdoor attack against diffusion models needs to poison half of the dataset [13]. Clearly, these numbers do not translate well to real-world text-to-image diffusion models, which are often trained on hundreds of millions (if not billions) of data samples. Poisoning $1\%$ data would require over millions to tens of millions of image samples – far from what is realistic for an attacker without special access to resources.
@@ 4 A Simple “Dirty-Label” Poisoning Attack
Note that this dirty-label attack involves attackers uploading images tagged with incorrect ALT-text. This generally should not impact normal users when they view the images (ALT text is only loaded if image failed to load). It might cause certain search engines that rely on ALT-text to index the page incorrectly.
@@ 6.1 Experimental Setup
Nightshade Attack Configuration.   Following the attack design in §5.3, we randomly select $5$K samples from LAION-5B (minus LAION-Aesthetic) as the natural dataset $\{\mathsf{Text/Image}\}$. We ensure they do not overlap with the 100K training samples in Table II. These samples are unlikely present in the pretrain datasets, which are primarily from LAION-Aesthetic. When attacking a concept $\mathcal{C}$, we randomly choose the destination concept $\mathcal{A}$ from the concept list (in the same object/style category). For guided perturbation, we first crop all image data into 512 x 512 squares (input size of diffusion models) and then we follow prior work to use LPIPS budget of $p=0.07$ and run an Adam optimizer for 500 steps [79, 14, 84]. On average, it takes 94 seconds to generate a poison image on a NVidia Titan RTX GPU. Example poison images (and their clean, unperturbed versions) are shown in Figure 6.
@@ 6.2 Attack Effectiveness
Visualizing changes in model internals.   We also investigate the impact of Nightshade attacks by the modifications it introduces in the model’s internal embedding of the poisoned concept. Specifically, we study the cross-attention layers, which encode the relationships between text tokens and a given image [85, 82]. Higher values are assigned to the image regions that are more related to the tokens, visualized by brighter colors in the cross-attention map. Figure 13 plots the cross-attention maps of a model before and after poisoning model (SD-V2 with $200$ poison data) for two object concepts targeted by Nightshade (“hat” and “handbag”). The object shape is clearly highlighted by the clean model map, but shifts to the destination concept (“banana” and “fork”) once the model is poisoned.
@@ 6.3 Impact of Clean Training Data
Subsequent continous training on clean data only.   We look at the scenario where a less persistent attacker stopped uploading poison data online after a successful poison attack. Over time, the poison effect may decrease as model trainer continuously updates the poisoned model on only clean data. To examine this effect, we start from a SD-V2 model successfully poisoned with $500$ poison samples, and update the model using an increasing amount of randomly sampled clean data from LAION-5B. Figure 25 in the Appendix shows that the attack success rate does decrease with the # of new clean data. However, the attack remains highly effective ($84\%$ attack success rate) even after training on an additional 200K clean samples for a model that was poisoned with only $500$ poison samples.
@@ 6.4 Bleed-through to Other Concepts
6.5 Composability Attacks
@@
Figure 17 illustrates the impact of these attacks with example images generated on prompts not targeted by any poison attacks. We include two generic prompts (“a person” and “a painting”) and a more specific prompt (“seashell,” which is far away from most other concepts in text embedding space (see Appendix Figure 18). Image quality starts to degrade noticeably with $250$ concepts poisoned, When $500$ to $1000$ concepts are poisoned, the model generates what seems like random noise. For a model training from scratch (LD-CC), similar levels of degradation requires $500$ concepts to be poisoned (Table IX in Appendix). The degradation on the entire model is likely because poison data (image & text) is “misaligned”, it increases the difficulty of learning text-image alignment in the model and corrupts the cross-attention layer. We leave further analysis of its cause to future work.
@@ 6.6 Attack Generalizability
Attack performance on diverse prompts.   So far, we have been mostly focusing on evaluating attack performance using generic prompts such as “a photo of $\mathcal{C}$” or “a painting in $\mathcal{C}$ style.” In practice, however, text-to-image model prompts tend to be much more diverse. Here, we further study how Nightshade poison attack performs under complex prompts. Given a poisoned concept $\mathcal{C}$, we follow prior work [37] to generate $4$ types of complex prompts (examples shown in Table VI). More details on the prompt construction can be found in Section 4 of [37]. We summarize our results in Table VI. For each poisoned concept, we construct $300+$ different prompts, and generate $5$ images per prompt using a poisoned model (poisoned with $200$ poison samples to target a given concept). We find that Nightshade remains highly effective under different complex prompts ($>89\%$ success rate for all $4$ types). In addition, we further show the attack remains successful on extremely long prompts in Appendix A.4.
@@ 7 Potential Defenses
Automated image captioning.   Next, we look at a defense method where model trainer completely removes the text prompt for all training data in order to remove the poison text. Once removed, model trainer can leverage existing image captioning tools [95, 96] to generate new text prompts for each training image. Similar approaches have been used to improve the data quality of poorly captioned images [97, 98].
@@
Gradient-based Outlier Detection.   Lastly, we look at whether an attacker can leverage outlier detection to identify poison images [99, 100, 46] through anomalies in their gradient. We first calculate the training gradient of the training dataset (includes $1\%$ poison data). Then we run one-class SVM detector on the gradient. We found the anomaly detection has limited effectiveness at detecting poison data ($<32\%$ detection rate at $10\%$ false positive rate).
@@ 8 Poison Attacks as Copyright Protection
We have released Nightshade as an independent app for Windows and Mac platforms. Response from the global artist community has been overwhelming, with 250K downloads in the first 5 days of release. Since then, we have began discussions with several companies in different creative industries who wish to deploy Nightshade on their copyrighted content. Finally, relevant model companies Google, Meta, Stability.ai and OpenAI have all been made aware of this work prior to this publication.
@@ 9 Conclusion
Acknowledgements
We thank our anonymous reviewers and shepherd for their insightful feedback. We also thank Karla Ortiz, Eva Toorenent, Katria Raden, Jingna Zhang, Kelly McKernan, Jon Lam, Sarah Andersen, Zakuga Mignon, Yujin Choo, Steven Zapata, Greg Rutkowski, Viktoria Sinner, Katharina Jahn, Paloma McClain, Jess Cheng, Erick Diego Castillo Chavez, and many other artists, without whom this project would not be possible. This work is supported in part by NSF grants CNS-2241303, CNS-1949650, and the DARPA GARD program. Opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of any funding agencies.
@@ A.3 Additional Results of Simple Dirty-Label Poisoning Attacks
All poison concepts used in the paper.   The following is all the concepts (from MSCOCO and WikiArt datasets) we used in the paper.
  • MSCOCO: ’shoe’, ’umbrella’, ’sink’, ’pizza’, ’airplane’, ’suitcase’, ’person’, ’sheep’, ’remote’, ’laptop’, ’surfboard’, ’racket’, ’spoon’, ’eye glasses’, ’desk’, ’street sign’, ’house’, ’hat’, ’tv’, ’pyramid’, ’frisbee’, ’knife’, ’fork’, ’clock’, ’microwave’, ’toothbrush’, ’mirror’, ’chair’, ’boat’, ’keyboard’, ’bicycle’, ’cow’, ’kite’, ’snowboard’, ’traffic light’, ’glove’, ’backpack’, ’mushroom’, ’sandwich’, ’cat’, ’scissors’, ’bird’, ’apple’, ’carrot’, ’panda’, ’motorcycle’, ’hot dog’, ’plate’, ’bus’, ’phone’, ’train’, ’bowl’, ’dog’, ’bench’, ’table’, ’toilet’, ’lawyer’, ’book’, ’window’, ’refrigerator’, ’elephant’, ’broccoli’, ’donut’, ’banana’, ’astronaut’, ’zebra’, ’vase’, ’bear’, ’truck’, ’fire hydrant’, ’whale’, ’skis’, ’handbag’, ’cake’, ’giraffe’, ’potted plant’, ’toaster’, ’castle’, ’tie’, ’blender’, ’bottle’, ’car’, ’skateboard’, ’door’, ’oven’, ’bed’, ’couch’, ’hair drier’, ’cup’, ’orange’, ’wine glass’, ’mouse’, ’horse’.
  • WikiArt: ’High Renaissance’, ’Ukiyo e’, ’Northern Renaissance’, ’Pointillism’, ’Symbolism’, ’Pop Art’, ’Romanticism’, ’Mannerism Late Renaissance’, ’Early Renaissance’, ’Baroque’, ’Action painting’, ’Fauvism’, ’Color Field Painting’, ’Minimalism’, ’Naive Art Primitivism’, ’New Realism’, ’Realism’, ’Post Impressionism’, ’Contemporary Realism’, ’Expressionism’, ’Synthetic Cubism’, ’Analytical Cubism’, ’Rococo’, ’Impressionism’, ’Art Nouveau Modern’, ’Cubism’, ’Abstract Expressionism’.
@@
Figure 25: Nightshade’s attack success rate (CLIP-based) decreases when model trainer continuously trains an already-poisoned model on an increasing number of clean data. The base model is poisoned with 100, 300, and 500 poison data samples.
A.4 Additional Results
Poison bleed-through.   We evaluate the “related” concept bleed-through effects between artists and the art styles they are known for. We include $195$ artists associated with $28$ styles from the Wikiart dataset [73]. We poison each art style $\mathcal{C}$, then test poison’s impact on generating painting of artists whose style belong to style $\mathcal{C}$, without mentioning the poisoned style $\mathcal{C}$ in the prompt, e.g., query with “a painting by Picasso” for models with “cubism” poisoned. Table VII shows that with $200$ poison data on art style, Nightshade achieves $>91\%$ CLIP attack success rate on artist names alone, similar to its performance on the poisoned art style.
Attack performance on complex prompts.   We also evaluate poison performance on longer and more complex prompts. We sample 20 prompts that contains 30+ words from Midjourney showcase website. We then manually replace the concept in the prompt to the poisoned concept (e.g., “front view of a hyper realistic [CONCEPT] with white and blue body …”). The poison attack remains successful for complex prompts ($>80\%$ CLIP attack success rate).
Appendix B Meta-Review
The following meta-review was prepared by the program committee for the 2024 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
B.1 Summary
The paper introduces Nightshade, a highly optimized data poisoning attack against diffusion-based text-to-image models, demonstrating the vulnerability of these models to prompt-specific poisoning attacks with a small number of optimized samples. The attack exploits concept sparsity and stealthily corrupts image generation for specific concepts, posing risks to model integrity and suggesting a tool for intellectual property protection.
B.2 Scientific Contributions
  • Identifies an Impactful Vulnerability
  • Provides a Valuable Step Forward in an Established Field
B.3 Reasons for Acceptance
  1. 1.
    This paper illustrates a significant vulnerability in state-of-the-art models to data poisoning. The vulnerability is present even when large training datasets are used.
  2. 2.
    The proposed attack, Nightshade, is efficient, requiring fewer than 100 poisoned samples to significantly influence model outputs.
  3. 3.
    The paper is well-written, with a clear presentation and comprehensive evaluation of the proposed attack and defense mechanisms, and the experimental results are comprehensive.